# New Aspects of Analyzing Amyloid Fibrils

### v1

@@ 4.2. The Vietoris-Rips (VR) complex
The Vietoris-Rips (VR) complex (denoted by $\operatorname{VR}(X)$) of $X$ with parameter $\epsilon$ is the set of all simplices $\sigma=\{x_{0},\cdots,x_{k}\}$, such that $d(x_{i},x_{j})\leq 2\epsilon$ for all $(i,j)$.
@@
(24) $\emptyset=\operatorname{VR}(X)_{0}\subseteq{\operatorname{VR}(X)_{1}}\subseteq{\cdots}\subseteq{\operatorname{VR}(X)_{m}}$
@@
Figure 5 is the PD of the C-terminal fragment of chain A of 7OB4 using the VR complex. The data set is given by taking the coordinates of $C_{\alpha}\text{s}$ on the C-terminal fragment. Since each chain of 7OB4 is almost planner, we only computed the persistence on $H_{0}$ and $H_{1}$. Red points and blue points represent the 0-dimensional and 1-dimensional homological generators respectively in the Fig.5.
Figure 5. The PD of the C-terminal fragment of 7OB4. Red points and blue points represent the 0-dimensional and 1-dimensional homological generators, respectively. The unit is the Angstrom (Å).
@@ 4.3. Applications to Analysis of the Structure of Amyloid fibrils
(26) $\text{RMSD}=\sqrt{\frac{\sum\limits_{i=1}^{n}(c_{i}-c_{i}^{\prime})^{2}}{n}}.$
It is clear that the RMSD only considers the difference of pairwise distance. Here, we adopted the PH to quantify the difference between two amyloid fibril structures from multiscales. The pseudo-code of comparing two amyloid fibrils is shown as the Algorithm 2. The results are listed in Table3.
Amyloid fibrils RMSD (Å) $W_{\infty}$(Å) $W_{2}$(Å) $W_{1}$(Å)
6SDZ-8ADE (N) 0.393 0.053 0.206 0.991
6SDZ-8ADE (C) 0.449 1.028 1.142 4.428
Table 3. Indicators to quantify the difference of the structure of two layers of amyloid fibrils (PDB entries: 6SDZ and 8ADE). The letter in parentheses represents the terminal fragment of chain A of each amyloid fibril.
We found that the RMSDs of two fragments are both less than 1Å. It indicates that the structure of 6SDZ and 8ADE are considered as almost the same in the RMSD context. However, the patient carries 6SDZ and the patient carries 8ADE have different clinical manifestations [66], which implies that there should be differences between the structure of these amyloid fibrils. We find that the global difference between these amyloid fibrils can be captured by the $W_{\infty}$ distance, while the local difference between these amyloid fibrils is scaled up as the exponent order $q$ decreases. PH considers the higher-order relations among these $C_{\alpha}\text{s}$, which could give us more information to distinguish the structure of amyloid fibrils.

### v3

@@ 4.2. The Vietoris-Rips (VR) complex
The Vietoris-Rips (VR) complex (denoted by $\operatorname{VR}(X)$) of $X$ with parameter $\epsilon$ is the set of all simplices $\sigma=\{x_{0},\cdots,x_{k}\}$, such that $d(x_{i},x_{j})\leq 2\epsilon$ for all $(i,j)$ [9].
@@
(26) $X=\operatorname{VR}(X)_{0}\subseteq{\operatorname{VR}(X)_{1}}\subseteq{\cdots}\subseteq{\operatorname{VR}(X)_{m}}$
@@
Figure 8 is the PD of the C-terminal fragment of chain A of V30MA-DIMER using the VR complex. The data set is given by taking the coordinates of $C_{\alpha}\text{s}$ on the C-terminal fragment. Since each chain of V30MA-DIMER is almost planner, we only computed the persistence on $H_{0}$ and $H_{1}$. Red points and blue points represent the 0-dimensional and 1-dimensional homological generators respectively in the Fig.8 and Fig.9A.
Figure 8. The PD of the C-terminal fragment of V30MA-DIMER. Red points and blue points represent the 0-dimensional and 1-dimensional homological generators, respectively. The unit is the Angstrom (Å).
@@ 4.3. Applications to Analysis of the Structure of Amyloid fibrils
(28) $\text{RMSD}=\sqrt{\frac{\sum\limits_{i=1}^{n}||c_{i}-c_{i}^{\prime}||^{2}}{n}}.$
It is clear that the RMSD only considers the difference of pairwise distance (Fig.9B). Here, we adopted the PH to quantify the difference between two amyloid fibril structures from multiscales (Fig.9A). The pseudo-code of comparing two amyloid fibrils is shown as the Algorithm 2. Please contact us if you require the source code. The results are displayed in the Fig.9C.
Figure 9. Dissimilarity measurements. (A) Flowchart of capturing dissimilaritues between the C-terminals on chain A of V30Ma and V30Ma-dimer by adopting TDA. (B) The alignment between chain A of ATTRs. (C) Dissimilarity measurements based on various methods. The capital letters N and C represent the terminal fragments of chain A of each ATTR, respectively.
We found that the RMSDs of two fragments are both less than 1Å. It indicates that the structure of ATTRs are considered as almost the same in the RMSD context. However, for example the patient carries 6SDZ and the patient carries 8ADE have different clinical manifestations [69], which implies that there should be differences between the structure of these amyloid fibrils. We find that the global difference between these amyloid fibrils can be captured by the $W_{\infty}$ distance, while the local difference between these amyloid fibrils is scaled up as the exponent order $q$ decreases. PH considers the higher-order relations among these $C_{\alpha}\text{s}$, which could give us more information to distinguish the structure of amyloid fibrils (Fig.9C).
@@ 5      $\text{PD}_{i}\leftarrow$ the persistent homology analysis of $\{K^{i}_{j}\}_{j=1}^{n}$
5. Discussion
In this study, we introduced geometrical and topological tools for analysing structures of amyloid fibrils. Geometrically, we defined the (truncated) $n$-hop distance to analyze the potential conformational transformations during the formation of ATTR fibrils from TTR tetramers. We found that the disruption of the $\alpha$-helix alters the interactions between the original secondary structures, which affected the values of hop distance significantly. Moreover, based on the special “pseudo-2D” structure of amyloid fibril, we defined a novel discrete curvature and torsion for structural analysis of intra- and inter- layers of ATTR. The torsions at the carbonyl carbons are significantly greater than those at other backbone atoms, and the absolute value of torsions is negatively correlated with the squared distance differences between the atoms on nearby layers that form the backbone hydrogen bonds. We concluded that the anomalies in the torsions of carbonyl carbons reflect slight effects of the layer-layer hydrogen bonds, the widest range of interactions in cross-$\beta$ structures. It is still unclear which specific physical properties of amyloid fibrils are relevant to the torsions. More datasets and experiments are needed to establish the link between these geometrical trends and macroscopic physical properties.
Additionally, we compared the differences between various polymorphs of ATTR adopting topological analysis. Compared to RMSD, TDA can more clearly distinguish the local differences between polymorphs and is expected to be useful for classifying different fibrillar polymorphs formed by the same amyloid protein. Unlike TTR polymorphs, which have low overall differences (small RMSDs), many amyloid proteins can form significantly different polymorphs in different contexts. However, the application of TDA to the analysis of these polymorphs is still under development due to the deficiency of structures of amyloid fibrils. Because of the small acount of data, the superiority of TDA has not clearly stand out compared with other techniques of data dimensionality reduction, such as PCA can not distinguish I84S-P1 and I84S-P2(Fig.10), but the others can be distinguished. Whereas, TDA has advantages that these methods do not, such as the multiscale characterization of structures of amyloid fibrils and the interpretative of the extracted structural features, we believe that the combination of TDA with machine learnings could give novel insights in the research of amyloid fibrils in the future.
Figure 10. PCA can not distinguish I84S-p1 and I84S-p2 well.